Circular No. 7539

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

S/2000 S 10 Another Saturnian satellite has been reported by the IAUC 7512 team:

2000 UT α_{2000} δ_{2000} R Observer
Sept. 23.47373 3^{h}58^{m}47.^{s}49 +18^{o}34′49′′.6 23.1 Gladman
24.48641 3 58 42.24 +18 34 36.7 23.2  _{”}
Nov. 24.20963 3 42 42.59 +17 51 40.7 23.0 Petit
27.18394 3 41 40.47 +17 48 53.4  Kavelaars

B. Gladman and J. Kavelaars (Mauna Kea). 3.6-m Canada-France-Hawaii

Telescope. Measurer Gladman. J.-M. Petit and B. Gladman (ESO). 2.2-m reflector. Measurer Gladman. J. Kavelaars and P. Nicholson (Palomar). 5-m refl. Measurer Kavelaars.

2000/01 α_{2000}  δ_{2000}  Δ  r_{S}  ϵ Δα′′  Δδ′′
Dec. 2 3^{h}40.^{m}02 +17^{o}44.^{′}5 8.037 0.153 166.^{o}7 962+ 2192+
12 3 36.79 +17 36.1 8.086 0.157 155.8 804+ 2219+
22 3 33.96 +17 29.1 8.165 0.161 144.9 647+ 2232+
Jan. 1 3 31.68 +17 24.1 8.269 0.164 134.2 496+ 2233+
11 3 30.07 +17 21.5 8.395 0.167 123.6 357+ 2226+

OUTER SATELLITES OF THE GIANT PLANETS During the past year and a half the number of known outer satellites

(or candidate outer satellites) of the giant planets has more than doubled. Although the Central Bureau is charged with providing provisional designations for new satellite candidates, to provide all the astrometric, orbital and ephemeris information requires a prohibitive amount of space on these Circulars. Readers should be aware that, by arrangement with IAU Commission 20, full astrometric and orbital information is now being included in the Minor Planet Circulars, and full details on S/2000 S 5–S 10 will appear in the Dec. 11 batch. In the future, the Minor Planet Electronic Circulars will be used to provide the initial information more completely.

Further observations are particularly required of S/2000 S 3, while recent observations of S/2000 S 4 indicate a significant ephemeris revision:

2000/01 α_{2000}  δ_{2000}  Δ  r_{S}  ϵ Δα′′  Δδ′′
Dec. 2 3^{h}40.^{m}43 +17^{o}19.^{′}3 8.200 0.072 166.^{o}6 1314+ 678+
12 3 37.60 +17 14.1 8.250 0.079 155.8 1501+ 898+
22 3 35.12 +17 10.2 8.326 0.086 145.1 1649+ 1096+
Jan. 1 3 33.15 +17 08.1 8.428 0.092 134.5 1761+ 1270+
11 3 31.80 +17 08.1 8.552 0.098 124.0 1844+ 1421+
2000 December 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Brian G. Marsden